The force of air resistance on a raindrop is
5.8 × 10−5 N when it falls with a terminal
velocity of 5.7 m/s.
The acceleration of gravity is 9.8 m/s
2
.
What is the mass of the raindrop?
1
Expert's answer
2019-11-21T08:50:26-0500

When a raindrop is moving with terminal velocity, the net force on it is zero, so


"mg=F_{\rm drag}"

Therefore


"m=\frac{F_{\rm drag}}{g}=\frac{5.8\times 10^{-5}\:\rm N}{9.8\:\rm N/kg}=5.92\times 10^{-6}\: \rm kg=5.92\:mg"